OpenStudy (anonymous):

Find the sum of the first 100 terms of the series in which tn = 2n - 1, n=1,2,3,...

hartnn (hartnn):

find the first term a1, put n=1 in an, what u get ?

hartnn (hartnn):

n=1 in tn*

hartnn (hartnn):

are u there @jufrando

OpenStudy (anonymous):

yeah sorry i was on another question

hartnn (hartnn):

so whats t1 ?

OpenStudy (anonymous):

so the question will tn= 2(1) -1? which is 1

OpenStudy (anonymous):

will be*

hartnn (hartnn):

yup, t1=1=1st term now find last term, t100, buy putting n=100

OpenStudy (anonymous):

thanks, so the answer is 199?

hartnn (hartnn):

thats last term = t100 = 199, answer is the sum given by S= n/2(1st term + last term)

hartnn (hartnn):

s= 100/2 (t1+t100)

OpenStudy (anonymous):

oh ok i get it now thank you, which equals to 10,000

hartnn (hartnn):

yup.